# Proxy Prompt: Endowing SAM & SAM 2 with Auto-Interactive-Prompt for Medical Segmentation

### v1

@@ Abstract
In this paper, we aim to address the unmet demand for automated prompting and enhanced human-model interactions of SAM and SAM2 for the sake of promoting their widespread clinical adoption. Specifically, we propose Proxy Prompt (PP), auto-generated by leveraging non-target data with a pre-annotated mask. We devise a novel 3-step context-selection strategy for adaptively selecting the most representative contextual information from non-target data via vision mamba and selective maps, empowering the guiding capability of non-target image-mask pairs for segmentation on target image/video data. To reinforce human-model interactions in PP, we further propose a contextual colorization module via a dual-reverse cross-attention to enhance interactions between target features and contextual-embedding with amplifying distinctive features of user-defined object(s). Via extensive evaluations, our method achieves state-of-the-art performance on four public datasets and yields comparable results with fully-trained models, even when trained with only 16 image masks.
WANG XINYI
The Hong Kong Polytechnic University
haylee-xinyi.wang@connect.polyu.hk
KANG HONGYU
The Hong Kong Polytechnic University
hong-yu.kang@connect.polyu.hk
WEI PEISHAN
The Hong Kong Polytechnic University
peishan.wei@polyu.edu.hk
SHUAI LI
The Hong Kong Polytechnic University
sshuai.li@connect.polyu.hk
YU SUN
The Hong Kong Polytechnic University
sunxinbrier@foxmail.com
SAI KIT LAM [^1] Sai Kit LAM and Yong-ping ZHENG are also affiliated with the Department of Biomedical Engineering and the Research Institute for Smart Ageing. Both authors are corresponding authors.
The Hong Kong Polytechnic University
saikit.lam@polyu.edu.hk
YONGPING ZHENG [^1] Sai Kit LAM and Yong-ping ZHENG are also affiliated with the Department of Biomedical Engineering and the Research Institute for Smart Ageing. Both authors are corresponding authors.
The Hong Kong Polytechnic University
yongping.zheng@polyu.edu.hk
@@ 1 Introduction
The core design of PPG lies in the novel Contextual Selective Module (CSM) and Contextual Colorization Module (CCM), which are dedicatedly configurated for automated prompting and high-level human-model interactions. First, CSM is introduced to enable adaptive selection of the most representative contextual information from “non-target” data for the “target” data, achieving cross-data guidance. Besides, CSM contains Vision Mamba, Bridge Unit, and Selective Map, implementing a 3-step selection process: (i) input-driven, (ii) object-guided, and (iii) target image/frame relevance selection to support both cross-video tasks and cross-image prompting.
Second, CCM is devised to reinforce human-model interaction, enabling the model to interpret diverse user needs as indicated by different masks (e.g. single/multiple objects). This aim is achieved by leveraging dual-reverse cross-attention to enhance the representation of contextual embedding. Finally, the PP, effectively capturing specific object features, is generated. Such PPG-based strategy presents a simple yet efficient architecture to enhance clinical-friendliness of SAM and SAM 2, even in few-shot settings. Furthermore, with all original parameters frozen, our design can function as a flexible, plug-and-play module and can continuously adapt to the ever-evolving LVMs beyond SAM and SAM 2.
@@ 3.1 Contextual Selective Module
We leverage Vision Mamba [33]’s selective input capabilities, establish communication between different objects, and design a filtering mechanism to refine the contextual embedding. As illustrated in Fig. 3, support pairs $\{{s}_{i}\}_{i=1}^{N}$ are initially processed by the Vision Mamba encoder to extract feature $\mathbf{V}$ with input-adaptive parameters, marking the first selective step. Secondly, within the designed Bridge Unit, we enable inter-object communication across the concatenated features, thereby selecting information from object-level to form $\mathbf{A}_{\text{agg}}$. Thirdly, features ($\mathbf{F}_{\text{x}}$) and ($\mathbf{F}_{\text{sup}}$) are used to compute a $\mathbf{Selective}$ map. Each row in this map representing a given patch from a support image will positively influence a specific patch in $\mathbf{X}$. This selective map subsequently filters the aggregated information $\mathbf{A}_{\text{agg}}$ to derive the most valuable information as the contextual embedding $\mathbf{E}_{\text{ctx}}$. The following sections will detail each component of this pipeline.
First Selection Step: Vision Mamba. We use the encoder of Vision Mamba (Vim)  [33] as the contextual encoder, due to its natural alignment with our needs through its input-dependent selection mechanism. This mechanism enables parameters (i.e., $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$ and $\mathbf{C}$) influencing interactions within the image to adapt as functions of the input, allowing features to be extracted based on the input itself. Specifically, Vim first flattens the input 2-D image into patches and form the token sequence $x$, which is progressively transformed into the output token sequence $y$ according to the following formula, ultimately being encoded by us into the feature matrix.
@@
where $\mathbf{F}_{\text{cat}}\in\mathbb{R}^{K\times(N\times(C^{v}+C))\times H\times W}$ and $\mathbf{A}_{agg}\in\mathbb{R}^{N\times C\times(K\times H\times W)}$ represents the aggregated features of the support set.
Third Selection Step: Selective Map. In order to select the key information for the target features, we compute the selective map based on $\mathbf{F}_{x}$ and $\mathbf{F}_{\text{sup}}$. First, $\mathbf{F}_{x}$ and $\mathbf{F}_{\text{sup}}$ are flattened into $\mathbf{\hat{F}}_{x}\in\mathbb{R}^{C\times(H\times W)}$ and $\mathbf{\hat{F}}_{\text{sup}}\in\mathbb{R}^{C\times(K\times H\times W)}$, respectively. Then, map is calculated through matrix multiplication, and the calculation is performed using the following equation:
@@ 3.2 Contextual Colorization Module
The CCM consists of four identical blocks, one of which is detailed in Fig. 3. In each block, $\mathbf{\hat{F}}_{\text{x}}$ and $\mathbf{E}_{\text{ctx}}$ are each passed through a learnable projection layer to reduce their dimensions to $\mathbf{\hat{F}}_{\text{x}}\in\mathbb{R}^{N\times C\times(H\times W/2)}$ and $\mathbf{E}_{\text{ctx}}\in\mathbb{R}^{N\times C\times(H\times W/2)}$, respectively. Subsequently, $\mathbf{E}_{\text{ctx}}$ is integrated into the target image features to guide the model in identifying specific regions expected to be “colored”. This integration occurs by adding cross-attention-processed information back into $\mathbf{\hat{F}}_{\text{x}}$, followed by a feed-forward layer (FFN) [23] for contextual reasoning, yielding $\mathbf{\hat{F}}_{\text{x\_next}}$, which then serves as the next target feature input for the next block. This process is represented as follows:
@@
After passing through the final block, the iteratively refined context embedding fully comprehends the user-defined segmentation intent from the support set, as well as the corresponding image features derived from the target image. This enhanced context embedding subsequently serves as $N$ prompts $\mathbf{P}$ for the SAM or SAM 2 model, where $\mathbf{P}\in\mathbb{R}^{N\times(H\times W/16)\times C}$.
@@ 3.3 Loss Function
Figure 4: Comparison on the video dataset using five different support pairs, showing the average results for two objects. Results show that our method achieves SOTA performance over existing models and demonstrates stability with a maximum fluctuation of only 1.0%.
Figure 5: Comparison on one representative video. The top left corner of each subplot displays the Dice scores for the segmentation of the public symphysis(PS_DSC) and fetal head (FH_DSC).
4 Experiment
4.1 Datasets
We conducted extensive experiments on four widely used publicly available datasets, including REFUGE2 [3], STARE [8], FH-PS-AOP(FPA) [11], and JNU-IFM [14]. Among these, STARE [8] contains only one single target, while the other datasets contain multiple segmentation targets. REFUGE2 [3], STARE [8] and FPA [11] are image datasets evaluated in the image segmentation task, whereas JNU-IFM [14] is a video dataset assessed in the video task. Further introduction to the datasets can be found in the supplementary materials.
4.2 Implementation Details
The “ViT-B” SAM model was used and initialized with pre-trained weights. Due to the limited training data, we freeze the pre-trained parameters of SAM to avoid overfitting. We applied LoRA [9] on original encoder and decoder for efficient fine-tuning on medical datasets. The patch size and embedding dimension of ViM [33] encoder were set to 16 and 192, respectively. Model optimization was performed using the Stochastic Gradient Descent (SGD) optimizer, with a momentum of 0.9, a learning rate of 0.01, and a weight decay of 0.0005. Data augmentation was applied to enrich the training process under limited data conditions. This included random variations in image brightness, rotation angles, and shear angles. Specifically, subtle adjustments were made to image saturation, brightness, and contrast within a range of 0.1 to simulate different settings and image gains. The maximum random rotation angle was set to 15 degrees, and the shear angle was limited to 10 degrees. For each training iteration, a consistent augmentation was applied to all images and masks in the support set.
4.3 Comparison with SOTA on Video Dataset
In this present work, we conducted a comprehensive comparison on the JNU-IFM dataset across four models: MedSAM 2, SAM 2-box, SAM 2-point, and Ours. MedSAM 2 and Ours utilize the image-mask pair to achieve automatic prompting in real-time scenarios, experiments settings can be found in the supplementary materials. Consequently, the simplest prompt in real-time is provided by Ours and MedSAM 2, followed by SAM-point with only one click, and finally SAM-box. Although mask prompts have superior performance based on the most complex information, they are practically overwhelming by a single operator to implement in real-time scenarios, thus serves as our “Upper”.
We conducted five experiments on the same test set across the four models to evaluate their performance and stability. We randomly excluded five videos from five patients and paired the first frame’s image and mask as five candidate support pairs. In each experiment, we altered different support pair as prompt to assess stability.
Method Training times Color fundus photography Ultrasound Average
@@ 4.3 Comparison with SOTA on Video Dataset
Med-SA (1-point) 3 85.6 83.0 45.4 71.3 81.4 73.3
Med-SA (box) 3 86.8 82.6 37.1 71.8 83.1 72.3
SAMed 3 86.9 84.3 15.8 66.3 87.0 68.1
AutoSAM 5 87.9 84.6 73.8 78.0 79.1 80.7
Ours 2 88.2 85.6 78.9 81.5 88.0 84.4
Table 1: Comparison of our model with SAM-based SOTAs in Dice Score (%), including train-free models (denoted as ${}^{\text{TF}}$), efficient fine-tuned models, and models trained on full data (denoted as $Upper^{*}$). ‘Training times’ records minimum number of trained models required to segment 5 objects, with dash(–) denoting inapplicability. Gray indicates the data were used in model pre-training. The $\bf best$ and $\underline{\text{second-best}}$ results (excluding Upper) are bolded and underlined, respectively, showcasing our SOTA performance among various models under same few-shot settings.
@@
The quantitative comparison results are shown in Fig. 4. From the averaged performance across five experiments, it can be observed that our method improves the dice score by 17.5% compared to MedSAM 2, which uses a similarly simple prompt strategy. Even when compared to SAM 2-box, which employs the most demanding prompt strategy, our method still achieves a 2.3% improvement and is comparable to the “Upper”. In terms of stability, MedSAM 2 is significantly affected by variations in the support pairs across the five trials. This results in a maximum dice score fluctuation of 29.1% for MedSAM 2, whereas our method maintains a far smaller fluctuation of only 1.0%.
The qualitative comparison results are shown in Fig. 5. Support pair 1 and 2 were chosen to demonstrate the stability of MedSAM 2 and our model under different prompts. Due to the blurred FH boundary in this case, both SAM 2-box and Ours (as well as the “Upper”) exhibited slight over-segment in FH segmentation. However, while SAM 2-box over-segmented the FH region to the left in $Frame_{t}$, our method constrains this expansion. Although MedSAM 2 showed a notable drop in PS segmentation performance with support pair 1, our method maintained consistent results.
@@ 4.4 Comparison with SOTA on Image Dataset
We extensively evaluated diverse comparison methods, including train-free foundation models, models efficiently fine-tuned on the specific-dataset, and traditional segmentation models. Experiment settings of prompts are in the supplementary materials. For train-free models such as SAM and MedSAM, we tested their performance with different prompt. Since MedSAM have been pre-trained on a large medical dataset, we evaluated its performance on both the pre-trained FPA dataset and the unseen REFUGE2 dataset. For models requiring training, including Med-SA, SAMed, AutoSAM, and ours, we trained each model on the same 16 images until full convergence to evaluate their performance in few-shot settings. For traditional segmentation models, we used dataset-specific SOTA methods, including BEAL [25] for REFUGE2, nnUnet [10] for STARE, and Segnet [2] for FPA. The results of the three models, as claimed in their papers, were obtained under full data training settings and thus were treated as the “Upper”.
@@
The quantitative comparison results are presented in Tab. 1. MedSAM and AutoSAM showed strong performance in both train-free and few-shot settings, probably due to pre-training on large medical datasets or object-specific training. However, MedSAM’s performance dropped significantly on unseen datasets and struggled with intricate structure such as vessel in the STARE dataset. AutoSAM, on the other hand, underperformed on ultrasound datasets, possibly due to the challenges such as ultrasound imaging-associated speckle noise. Furthermore, AutoSAM requires training five models for five objects, whereas our method only requires training two models for two modalities with human-model interaction. This reduction in retraining enhances practicality in real-world clinics, where tasks are more diverse. Our method achieved SOTA performance under limited conditions, even comparable with traditional SOTAs trained on the full dataset.
The qualitative comparison results are presented in Fig. 6. For the REFUGE2 dataset, SAMed and AutoSAM performed relatively well, nevertheless both showed issues with missing regions in the boundaries for certain objects. On the vascular dataset STARE, MedSAM with box prompt exhibited ambiguity, confusing the objects to be segmented. For PS in the FPA dataset, most models tended to over-segment, with SAM-point, SAM-box, Med-SA, and AutoSAM extending the PS segmentation beyond the object’s lower right boundary, while SAM-everything and SAMed less segment in the lower left edge. For FH, most models tended to over-segment, while SAM-box and SAMed exhibited less segmentation at the upper left edge. This may be caused by blurring of the object boundary due to presence of speckle noise in the ultrasound dataset. Across all datasets, our method consistently produced superior and stable results.

### v2

@@ Abstract
In this paper, we aim to address the unmet demand for automated prompting and enhanced human-model interactions of SAM and SAM2 for the sake of promoting their widespread clinical adoption. Specifically, we propose Proxy Prompt (PP), auto-generated by leveraging non-target data with a pre-annotated mask. We devise a novel 3-step context-selection strategy for adaptively selecting the most representative contextual information from non-target data via vision mamba and selective maps, empowering the guiding capability of non-target image-mask pairs for segmentation on target image/video data. To reinforce human-model interactions in PP, we further propose a contextual colorization module via a dual-reverse cross-attention to enhance interactions between target features and contextual-embedding with amplifying distinctive features of user-defined object(s). Via extensive evaluations, our method achieves state-of-the-art performance on five public datasets and yields comparable results with fully-trained models, even when trained with only 16 image masks.
@@ 1 Introduction
The core design of PPG lies in the novel Contextual Selective Module (CSM) and Contextual Colorization Module (CCM), which are dedicatedly configurated for automated prompting and high-level human-model interactions. First, CSM is introduced to enable adaptive selection of the most representative contextual information from “non-target” data for the “target” data, achieving cross-data guidance. Besides, CSM contains Vision Mamba, Bridge Unit, and Selective Map, implementing a 3-step selection process: (i) input-driven, (ii) object-guided, and (iii) target image/frame relevance selection to support both cross-video tasks and cross-image prompting. Second, CCM is devised to reinforce human-model interaction, enabling the model to interpret diverse user needs as indicated by different masks (e.g. single/multiple objects). This aim is achieved by leveraging dual-reverse cross-attention to enhance the representation of contextual embedding. Finally, the PP, effectively capturing specific object features, is generated. Such PPG-based strategy presents a simple yet efficient architecture to enhance clinical-friendliness of SAM and SAM 2, even in few-shot settings. Furthermore, with all original parameters frozen, our design can function as a flexible, plug-and-play module and can continuously adapt to the ever-evolving LVMs beyond SAM and SAM 2.
@@ 3.1 Contextual Selective Module
We leverage Vision Mamba [44]’s selective input capabilities, establish communication between different objects, and design a filtering mechanism to refine the contextual embedding. As illustrated in Fig. 3, support pairs $\{{s}_{i}\}_{i=1}^{N}$ are initially processed by the Vision Mamba encoder to extract feature $\mathbf{V}$ with input-adaptive parameters, marking the first selective step. Secondly, within the designed Bridge Unit, we enable inter-object communication across the concatenated features, thereby selectively extracting object-level feature $\mathbf{A}_{\text{agg}}$. Thirdly, features ($\mathbf{F}_{\text{x}}$) and ($\mathbf{F}_{\text{sup}}$) are used to compute a $\mathbf{Selective}$ map. Each row in this map representing a given patch from a support image will positively influence a specific patch in $\mathbf{X}$. This selective map subsequently filters the aggregated information $\mathbf{A}_{\text{agg}}$ to derive the most valuable information as the contextual embedding $\mathbf{E}_{\text{ctx}}$. The following sections will detail each component of this pipeline.
First Selection Step: Vision Mamba. We use the encoder of Vision Mamba (Vim)  [44] as the contextual encoder, as its input-dependent selection mechanism is well-suited for learning similar modality features, thereby enhancing prompt effectiveness. This mechanism enables parameters (i.e., $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$ and $\mathbf{C}$) influencing interactions within the image to adapt as functions of the input, allowing features to be extracted based on the input itself. Specifically, Vim first flattens the input 2-D image into patches and form the token sequence $x$, which is progressively transformed into the output token sequence $y$ according to the following formula, ultimately being encoded by us into the feature matrix.
@@
where $\mathbf{F}_{\text{cat}}\in\mathbb{R}^{K\times N\times(C^{v}+C)\times H\times W}$ and $\mathbf{A}_{agg}\in\mathbb{R}^{N\times C\times(K\times H\times W)}$ represents the aggregated features of the support set.
Third Selection Step: Selective Map. In order to select the key information for the target features, we compute the selective map based on $\mathbf{F}_{x}$ and $\mathbf{F}_{\text{sup}}$. First, $\mathbf{F}_{x}$ and $\mathbf{F}_{\text{sup}}$ are flattened into $\mathbf{\hat{F}}_{x}\in\mathbb{R}^{C\times(H\times W)}$ and $\mathbf{\hat{F}}_{\text{sup}}\in\mathbb{R}^{C\times(K\times H\times W)}$, respectively. Then, the map is calculated through matrix multiplication, following  [5] and using the following equation:
@@ 3.2 Contextual Colorization Module
The CCM consists of four identical blocks, one of which is detailed in Fig. 3. In each block, $\mathbf{\hat{F}}_{\text{x}}$ and $\mathbf{E}_{\text{ctx}}$ are each passed through a learnable projection layer to reduce their dimensions to $\mathbf{\hat{F}}_{\text{x}}\in\mathbb{R}^{N\times C\times((H\times W)/2)}$ and $\mathbf{E}_{\text{ctx}}\in\mathbb{R}^{N\times C\times((H\times W)/2)}$, respectively. Subsequently, $\mathbf{E}_{\text{ctx}}$ is integrated into the target image features to guide the model in identifying specific regions expected to be “colored”. This integration occurs by adding cross-attention-processed information back into $\mathbf{\hat{F}}_{\text{x}}$, followed by a feed-forward layer (FFN) [30] for contextual reasoning, yielding $\mathbf{\hat{F}}_{\text{x\_next}}$, which then serves as the next target feature input for the next block. This process is represented as follows:
@@
After passing through the final block, the iteratively refined context embedding fully comprehends the user-defined segmentation intent from the support set, as well as the corresponding image features derived from the target image. This enhanced context embedding subsequently serves as $N$ prompts $\mathbf{P}$ for the SAM or SAM 2 model, where $\mathbf{P}\in\mathbb{R}^{N\times((H\times W)/16)\times C}$.
@@ 3.3 Loss Function
Method Training Images Color Fundus Photography Ultrasound Avg. Dice $\uparrow$ (%)
@@
Med-SA (1-point) 16 85.6 83.0 45.4 71.3 81.4 73.3
Med-SA (box) 16 86.8 82.6 37.1 71.8 83.1 72.3
SAMed 16 86.9 84.3 15.8 66.3 87.0 68.1
AutoSAM 16 87.9 84.6 73.8 78.0 79.1 80.7
Ours 16 88.0 85.5 81.6 85.3 89.3 85.9
Table 1: Comparison of our model with SAM-based SOTAs in Dice Score (%) on image datasets, including train-free models (denoted as ${}^{\text{TF}}$), efficient fine-tuned models, and models trained on full data (denoted as $Upper^{*}$). ‘Training Images’ denotes the average number of images used for training to segment five objects, with dash(–) denoting inapplicability. Gray indicates the data were used in model pre-training. The $\bf best$ and $\underline{\text{second-best}}$ results (excluding $Upper$) are bolded and underlined, respectively, showcasing our SOTA performance among various models under same few-shot settings.
Method Training data Avg. Dice $\uparrow$ (%)
With Label W/O. Label
$Upper^{*}$ 1134 (100%) 0 (0%) 88.1
DTC 58 (5.1%) 1076 (94.9%) 63.4
MLB-Seg 58 (5.1%) 1076 (94.9%) 78.3
nnUnet 58 (5.1%) 0 (0%) 84.1
Med-SA 58 (5.1%) 0 (0%) 76.0
SAMed 58 (5.1%) 0 (0%) 85.9
AutoSAM 58 (5.1%) 0 (0%) 70.1
Ours 58 (5.1%) 0 (0%) 87.4
Table 2: Comparison of our model against various SOTAs on MRI dataset, including semi-supervised models [20, 34], SAM-based models, traditional segmentation models, and 3D DSD-FCN [31] trained on the full dataset (denoted as $Upper^{*}$). ‘Training Data’ indicates the number of slices used for training and their percentage of the total dataset.
4 Experiment
4.1 Datasets and Implementation Details
We conducted extensive experiments on five widely used publicly available datasets, including REFUGE2 [6], STARE [11], FH-PS-AOP(FPA) [14], PROMISE12 [18] (3D MRI) and JNU-IFM [19] (Video). Among these, STARE [11] and PROMISE12 [18] contain only one single object, while the other datasets contain multiple segmentation objects. Further introduction in the supplementary.
Method Prompt Preload
Capable Avg.$\uparrow$
(%) Std.$\downarrow$
[…]
$Upper$ Mask 89.6 - -
SAM2 Box 78.6 - -
SAM2 Point 25.1 - -
MedSAM2 Mask 63.3 10.9 29.1
Ours Mask 80.9 0.3 1.0
Table 3: Comparison of our model with SAM2-based SOTAs on the video dataset, reporting Average Dice, standard deviation Dice, and max-min Dice difference. $Preload~{}Capable$ indicates whether the prompt can be pre-prepared for real-time use. Background shading from light to dark green represents increasing real-time usability, while SAM2-Mask, being nearly impractical, is treated as $Upper$ and marked in Gray. The $\bf best$ and $\underline{\text{second-best}}$ results (excluding $Upper$) are bolded and underlined, respectively.
Due to the limited training data, we freeze the pre-trained parameters of SAM to avoid overfitting. We applied LoRA [12] to the original encoder and decoder for efficient fine-tuning on medical datasets. Model optimization was performed using the Stochastic Gradient Descent (SGD) optimizer, with a momentum of 0.9, a learning rate of 0.01, and a weight decay of 0.0005. For image and video datasets, we randomly selected 16 patient-level images or videos for training, with the remaining samples used for testing. For 3D data, 58 labeled slices from 3 patients were used for training, while data from 10 additional patients were reserved for testing. Further details in the supplementary.
@@ 4.2 Comparison with SOTA on Image Dataset
We extensively evaluated diverse comparison methods, including train-free foundation models, models efficiently fine-tuned on the specific-dataset, and traditional models ($Upper$). Details are in the supplementary. MedSAM, pretrained on a large medical dataset, was evaluated on both its seen FPA dataset and the unseen REFUGE2 dataset. For trainable models (Med-SA, SAMed, AutoSAM, and ours), we trained each on the same 16 images until convergence.
@@
As shown in Tab. 1, MedSAM and AutoSAM showed strong performance in both train-free and few-shot settings, probably due to pre-training on large medical datasets or object-specific training. However, MedSAM’s performance dropped significantly on unseen datasets and struggled with intricate structure such as vessel in the STARE dataset. AutoSAM, on the other hand, underperformed on ultrasound datasets, possibly due to the challenges such as ultrasound imaging-associated speckle noise. Our method achieved SOTA performance under limited conditions, even comparable with traditional SOTAs trained on the full dataset. Fig. 4 shows that while SAMed and AutoSAM performed relatively well on REFUGE2 and FPA, both still struggled with boundary omissions. On STARE, MedSAM’s box prompt led to segmentation ambiguity. Across all datasets, our method delivered consistently superior and stable results.
4.3 Comparison with SOTA on MRI Dataset
Although not specifically designed for 3D data, our model shows strong potential. We evaluated it on the 3D MRI PROMISE12 dataset [18] (details in the supplementary). As shown in Tab. 2, it achieved the highest few-shot performance (87.4%), surpassing all SAM-based models and the traditional nnUnet under the same conditions. Notably, despite not leveraging any additional unlabeled data, our method surpassed the best semi-supervised model (MLB-Seg [34], 78.3%). Furthermore, we achieved results comparable to the fully-supervised $Upper^{*}$ (88.1%), with only a 0.7% gap, showcasing its promise for 3D segmentation.
4.4 Comparison with SOTA on Video Dataset
We conducted a comprehensive comparison on the JNU-IFM dataset across four models: MedSAM2, SAM2-box, SAM2-point, and Ours. Since Ours and MedSAM2 allow pre-prepared prompts from non-target data, they offer the simplest real-time usability, followed by SAM2-point (one-click), and lastly SAM2-box. Although mask prompts provide the most detailed information, they are impractical for a single operator in real-time and thus serve as our $Upper$. To assess the stability of MedSAM2 and Ours, we conducted five trials on the same test set, each using a different support pair randomly drawn from five pre-excluded videos. Further experimental details are in supplementary. As shown in Sec. 4.1, our method achieves the highest average Dice, improving by 17.6% over MedSAM2, which uses a simple prompt strategy, and by 2.3% over SAM2-box, which requires the most demanding prompt strategy. Further, our method yields comparable performance to the $Upper$, and presents excellent stability indicated by 0.3% Std. Dice (far greater than MedSAM2 of 10.9% Std. Dice and 29.1% fluctuation in Dice).
CSM CCM Avg. Dice $\uparrow$ (%) std. Dice $\downarrow$ (%)
85.9 2.6
85.2 3.1
84.3 3.0
84.0 3.1
Table 4: Ablation experiments on different PPG modules.
Figure 5: Selective Map visualization on support images with different mean square errors (MSE) to the target image.